1.   The maintenance cost for equipment has been $10,000 per year for the past six years. The interest rate was 6% per year compounding monthly for the first two years, 7% for the third year, and 8% per year compounding quarterly for the last three years. What is the equivalent maintenance cost at now?
Answer #1

The formula is as follows: where r = rate of interest p.a., n = number of times it compounds in a year (so, if monthly, then n=12), t = number of years.

So, first two years: Amount = 10000(1 + 0.005)^(24) = 11271.59

Third year: 11271.59(1+0.07) = 12060.6

last 3 years: 12060.6[(1+0.02)^12] = 15295.75. This is the final rate.